OpenStudy (anonymous):

Tuan simplified the following expression: (b7)(b-3)/b2 (The 7,-3, and 2 are exponents.) His final answer was in the form bn. If he simplified the expression correctly, what is the value of n, the exponent on b?

OpenStudy (anonymous):

When multiplying powers of the same base add the exponents; when dividing powers of the same base subtract the exponents.

OpenStudy (zugzwang):

Also keep in mind: \[\frac{a + b}{c}=\frac{a}{c}+\frac{b}{c}\]

OpenStudy (anonymous):

How is that relevant here, @zugzwang ?

OpenStudy (zugzwang):

It isn't. I'm sorry, I forgot that -3 was also an exponent MY BAD

OpenStudy (anonymous):

Ok, I thought there was something I was missing. Thank you for the clarification.